# Demystifying Spatial Confounding

### v2

@@ 4.4 Scenario 4: Dependence on non-spatial information
We now study the effect on the bias of non-spatial information in $\mathbf{x}$ by increasing the parameter $\sigma_{x}$. As the coefficient $\xi_{\mathrm{ns}}^{x}$ appears only in the denominator of the bias in both the spatial and the non-spatial model, we expect bias in both models to reduce. Specifically, we consider $\sigma_{x}\in\{0.3,1,3\}$. Other parameters remain constant at $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(0,0.5)$.
Figure 1 (bottom, right) shows that, as expected, the bias decreases for both the non-spatial and spatial model when $\sigma_{x}$ is increased. As before, the MSE of fitted values is always larger for the non-spatial model. In Section 3.3, we saw that non-spatial bias tends to exceed that of the spatial model, but if the covariate has little non-spatial information and confounding takes place at high frequencies, then it may be the other way around. This behaviour is observed in Figure 1 where the non-spatial bias is higher than the spatial, except in the case $\sigma_{x}=0.3$, i.e where $\mathbf{x}$ has the smallest amount of non-spatial information.
@@ 4.5 Scenario 5: Dependence on the smoothing parameter
Figure 2 shows that for $\sigma_{x}=0.3$, the spatial bias increases from 0, crosses that of the non-spatial model and stays above it, until they meet again for large $\lambda$. As expected, increasing non-spatial information in $\mathbf{x}$ decreases bias in general, and for $\sigma_{x}=1$ and $\sigma_{x}=2$, the bias of the spatial model stays below that of the non-spatial model. For each data generating process in Figure 2, we also show the median estimated smoothing parameter $\hat{\lambda}$ using the generalised cross-validation criterion. At the estimated values of $\lambda$ the bias is 0.4037, 0.1804 and 0.0362, from left to right. For $\sigma_{x}=0.3$, the generalised cross-validation estimate is in a relatively flat area of the curve where the spatial bias exceeds the non-spatial. For both of the other values of $\sigma_{x}$, although the spatial bias is bounded by the non-spatial bias, the estimate $\hat{\lambda}$ is at the steep part of the curve, which means that small changes in $\hat{\lambda}$ can lead to large changes in the bias.
@@ 5 Bias adjustment
As the expressions in Section 3 involve the unknown confounder $\mathbf{z}$, in practical applications, we cannot simply compute the bias to assess the impact of confounding. As illustrated in Section 3.3 and the Scenario 3 simulations above, we also cannot use the difference between the estimates in the spatial and non-spatial models as an indicator for the existence or otherwise of bias. Here, we propose a general framework for bias assessment and adjustment based on our results. That is, for a given data set and choice of spatial analysis model, we propose methods for estimating the bias under different assumptions for the data generation scenario. One assumption that is not usually highlighted in the spatial confounding literature is whether the covariate $\mathbf{x}$ has non-spatial information. However, as illustrated in Sections 3 and 4, the bias may exhibit fundamentally different behaviour depending on whether or not this assumption holds. As outlined below, this assumption is also important in the context of bias adjustment. hile the work on spatial scales is well-known, the results may be somewhat counter-intuitive, and there has been a tendency to worry more about confounding at low spatial frequencies despite this being the less problematic scenario. This is perhaps because low frequencies are often seen as more likely to be confounded in practice, with higher frequencies expected to capture the more characteristic behaviour of the variables. Moreover, as seen from our comparison between the spatial vs. the non-spatial model, low frequency confounding can also lead to large differences in the estimates in these two models. Historically, such differences were interpreted as a sign of confounding with the non-spatial estimate mistakenly assumed to be unbiased. Our results confirm that bias in the spatial model is unlikely to be a problem when confounding is at low frequencies, as long as there is (sufficient) non-spatial or unconfounded high frequency information in the covariate.
@@ 5.1 Case 1: $\mathbf{x}$ has non-spatial information
Spatial+ (Dupont et al.,, 2022) is a method for bias adjustment which utilises the fact that the non-spatial component $\mathbf{x}_{\mathrm{ns}}=\mathbf{U}_{\mathrm{ns}}\bm{\xi}^{x}_{\mathrm{ns}}$ of the covariate is unconfounded and can therefore be used to identify the covariate effect. Although originally introduced for models in which the spatial effect is a thin plate spline, the method can be formulated for any spatial analysis model as follows. The model (3) with no covariates is initially fitted to $\mathbf{x}$ to give the decomposition $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}+\mathbf{r}^{x}$ where $\mathbf{x}_{\mathrm{sp}}$ are the fitted values (estimating the spatial part of $\mathbf{x}$) and $\mathbf{r}^{x}$ the residuals. The covariate effect $\beta$ is then estimated by fitting the model (3) but with $\mathbf{x}$ replaced by $\mathbf{r}^{x}$. From Proposition 1, the bias in spatial+ is therefore $\langle\mathbf{r}^{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{x},\mathbf{r}^{x}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$. As the residuals $\mathbf{r}^{x}$ estimate the non-spatial component $\mathbf{x}_{\mathrm{ns}}$ which is orthogonal to $\mathbf{z}$, this shows why we can expect the bias to be eliminated. Indeed, even though $\mathbf{r}^{x}$ may contain a small proportion of potentially confounded spatial components, its non-spatial part is likely to dominate so that bias is negligible (as seen in the Scenario 4 simulations).
@@ 5.2 Case 2: $\mathbf{x}$ is fully spatial
Sometimes $\mathbf{x}$ has no non-spatial components, i.e. $\bm{\xi}^{x}_{\mathrm{ns}}=\mathbf{0}$. For example, $\mathbf{x}$ could be entirely generated from a smooth spatial process, or the dimension $p$ of the spatial effect in (3) could be as large as the sample size $n$ (such as in a regional spatial model with only one data measurement per region). The covariate vector is then contained in the space spanned by the spatial basis vectors and model (3) is unidentifiable. Here, the smoothing penalty not only improves the fit but acts as a regulariser. The model cannot be fitted without smoothing, and the apportionment of the covariate effect between the covariate and spatial terms of the model is directly related to how the smoothing is applied. Our expression for the bias still holds, however, unlike Case 1 above where the non-spatial component $\mathbf{x}_{\mathrm{ns}}$ is known to be unconfounded, we have no information of the confounding scenario and methods like spatial+ may not be effective. The behaviour of the bias will also generally be harder to predict. In particular, even if confounding is at low frequencies (making the numerator of the bias small), the bias could still be significant as the absence of non-spatial components in $\mathbf{x}$ means that the denominator could also potentially be very small, especially when the covariate is dominated by low frequency behaviours.
@@
As shown in Sections 3 and 4, spatial frequencies play a key role for the bias, and techniques such as Fourier and wavelet analysis are useful for understanding the behaviours of the observed variables $\mathbf{y}$ and $\mathbf{x}$ (as suggested, for example, in Keller and Szpiro, (2020)). However, our expressions show that it is the frequencies shared by $\mathbf{x}$ and $\mathbf{z}$ that cause bias, and the data has no information on how such frequencies are split between the two variables. Therefore, without further (untestable) assumptions for identifiability, we cannot estimate the bias. In the case where $\mathbf{x}$ and $\mathbf{z}$ are realisations of two stationary Gaussian spatial processes, Guan et al., (2023) use a Fourier transform to describe spatial frequencies and show that an assumption of unconfoundedness at high frequencies can be used to obtain identifiability. As noted in Guan et al., (2023), this assumption seems natural as the highest frequency spatial variation in a variable is more likely to be characteristic to that particular variable and therefore not confounded with other spatial variables. Corollary 1 shows that this idea extends to our framework in which spatial frequencies are defined directly in terms of the spatial analysis model (3) as described in Section 2.2. Unconfoundedness at the highest frequencies then corresponds to the assumption that $\xi^{z}_{\mathrm{sp},i}=0$ for frequencies above a certain threshold. Thus, under this assumption, the frequency components of $\mathbf{x}$ above the threshold can be used to obtain identifiability.
@@ 5.3 Capped spatial+
Based on our analysis in Section 5.2, we present an adapted version of spatial+ which can be used to assess the bias in the more challenging Case 2 scenario. As previously noted, unconfounded high frequencies in $\mathbf{x}$ have a similar effect on the bias as non-spatial information. Therefore, assuming the $k$ highest frequencies are unconfounded, a natural way to adapt spatial+ is to (i) decompose the covariate as $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}^{k}+\mathbf{r}^{k}$ where $\mathbf{x}_{\mathrm{sp}}^{k}=\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ and $\mathbf{r}^{k}=\sum_{i=n-k+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ with $\mathbf{u}_{\mathrm{sp},i}$ the $i$th spatial eigenvector, and (ii) replace $\mathbf{x}$ by the unconfounded component $\mathbf{r}^{k}$ in model (3). Proposition 1 shows that the bias of the corresponding covariate effect estimate is $\langle\mathbf{r}^{k},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{k},\mathbf{r}^{k}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$ which should therefore, in theory, be 00 due to the assumption of unconfoundedness. However, as the component $\mathbf{r}^{k}$ can also be modelled as part of the highest frequencies of the spatial effect in (3), the estimate of $\beta$ obtained in this way is highly sensitive to inaccuracies at these frequencies of the estimated spatial effect. Such inaccuracies easily arise in practice as the estimation is based on overall spatial smoothing, driven by the fitted values rather than the behaviour at individual spatial frequencies. In order to avoid such sensitivity, we adapt the model matrix of the regression in (ii) by removing the columns corresponding to the frequencies above $k$. This more explicitly reflects the assumption of unconfoundedness at these frequencies but allows the spatial effect to continue to adjust for confounders at frequencies below the cap. We call the estimate of $\beta$ obtained from (i) and the adapted (ii) the capped spatial+ estimate. By construction, if the assumption of unconfoundedness above the cap is true, the bias for this estimate is 00. In practice, the true value of the cap $k$ is unlikely to be known, and it may therefore be mis-specified at a value $k^{\prime}\neq k$. If $k^{\prime}<k$, $\mathbf{r}^{k^{\prime}}=\sum_{i=n-k^{\prime}+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ is simply a smaller component of $\mathbf{r}^{k}$ which is still orthogonal to the spatial part of the adjusted spatial model. Therefore, the bias should still be eliminated. However, if $k^{\prime}>k$, then $\mathbf{r}^{k^{\prime}}$ contains a potentially confounded component of $\mathbf{x}$, namely, the $k^{\prime}-k$ highest frequencies of $\mathbf{x}_{\mathrm{sp}}^{k}$, and by Corollary 1 the bias becomes $\sum_{i=n-k^{\prime}+1}^{n-k}\xi^{x}_{\mathrm{sp},i}\xi^{z}_{\mathrm{sp},i}w_{i}/\sum_{i=n-k^{\prime}+1}^{n}(\xi^{x}_{\mathrm{sp},i})^{2}w_{i}$. For comparison, the bias in the original spatial model (3) in this scenario is $\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\xi^{z}_{\mathrm{sp},i}w_{i}/\sum_{i=1}^{n}(\xi^{x}_{\mathrm{sp},i})^{2}w_{i}$. Thus, the numerator for the spatial model includes additional confounded frequencies contributing to the bias, but at the same time the denominator is also larger. Therefore, in this case, the capped spatial+ estimate may not reduce and could even increase the size of the bias of the spatial model.
@@
Finally, we note that the assumption of unconfoundedness does not have to be limited to the highest frequencies of the spatial effect. If we believe the unconfounded behaviour of $\mathbf{x}$ is in a different frequency range, we can easily adapt the capped spatial+ method by replacing the component $\mathbf{r}^{k}$ by the part of $\mathbf{x}$ that is believed to be unconfounded.

### v3

@@ 4.4 Scenario 4: Dependence on non-spatial information
We now study the effect on the bias of non-spatial information in the covariate $\mathbf{x}$ by increasing the parameter $\sigma_{x}$. As the coefficient $\xi_{\mathrm{ns}}^{x}$ appears only in the denominator of the bias in both the spatial and non-spatial models, we expect bias in both models to reduce. Specifically, we consider $\sigma_{x}\in\{0.3,1,3\}$. Other parameters remain constant at $(\xi_{\mathrm{sp,l}}^{x},\xi_{\mathrm{sp,h}}^{x})=(1,1)$, $(\xi_{\mathrm{sp,l}}^{z},\xi_{\mathrm{sp,h}}^{z})=(0,0.5)$.
Figure 1 (bottom, right) shows that, as expected, the bias decreases for both models when $\sigma_{x}$ is increased, and the case $\sigma_{x}=3$ illustrates that, for a sufficiently large non-spatial component of $\mathbf{x}$, the bias in the spatial model becomes negligible. As before, the MSE of fitted values is always larger for the non-spatial model. In Section 3.3, we saw that bias in the non-spatial model tends to exceed that of the spatial model, but if the covariate has little non-spatial information and confounding takes place at high frequencies, then it may be the other way around. This behaviour is observed in Figure 1 where the bias in the non-spatial model is higher than that of the spatial model, except in the case $\sigma_{x}=0.3$, i.e where $\mathbf{x}$ has the smallest amount of non-spatial information.
@@ 4.5 Scenario 5: Dependence on the smoothing parameter
Figure 2 shows that for $\sigma_{x}=0.3$, the spatial bias increases from 0, crosses that of the non-spatial model and stays above it, until they meet again for large $\lambda$. As expected, increasing non-spatial information in $\mathbf{x}$ decreases bias in general, and for $\sigma_{x}=1$ and $\sigma_{x}=2$, the bias of the spatial model stays below that of the non-spatial model. For each data generating process in Figure 2, we also show the median estimated smoothing parameter $\hat{\lambda}$ using the generalised cross-validation criterion. At the estimated values of $\lambda$ the bias is 0.4037, 0.1804 and 0.0362, from left to right. For $\sigma_{x}=0.3$, the generalised cross-validation estimate is in a relatively flat area of the curve where the bias in the spatial model exceeds that of the non-spatial model. For both of the other values of $\sigma_{x}$, although the spatial model bias is bounded by the non-spatial model bias, the estimate $\hat{\lambda}$ is at the steep part of the curve, which means that small changes in $\hat{\lambda}$ can lead to large changes in the bias.
@@ 5 Bias adjustment
As the expressions in Section 3 involve the unknown confounder $\mathbf{z}$, in practical applications, we cannot simply compute the bias to assess the impact of confounding. Moreover, as illustrated in Section 3.3 and the Scenario 3 simulations above, the difference between the estimates in the spatial and non-spatial models is a poor diagnostic for the existence or otherwise of bias. Here, instead we propose a general framework for bias assessment and adjustment based on our results. That is, for a given data set and choice of spatial analysis model, we propose methods for estimating the bias under different assumptions for the data generation scenario. One assumption that is not usually highlighted in the spatial confounding literature is whether the covariate $\mathbf{x}$ has non-spatial information. However, as illustrated in Sections 3 and 4, the bias may exhibit fundamentally different behaviour depending on whether or not this assumption holds. As outlined below, this assumption is also important in the context of bias adjustment.
@@ 5.1 Case 1: $\mathbf{x}$ has non-spatial information
Spatial+ (Dupont et al.,, 2022) is a method for bias adjustment which utilises the fact that the non-spatial component $\mathbf{x}_{\mathrm{ns}}=\mathbf{U}_{\mathrm{ns}}\bm{\xi}^{x}_{\mathrm{ns}}$ of the covariate is unconfounded and can therefore be used to identify the covariate effect. Although originally introduced for models in which the spatial effect is a thin plate spline, the method can be formulated for any spatial analysis model as the following two-step procedure.
Fit the model (3) with no covariates to $\mathbf{x}$ to obtain the decomposition $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}+\mathbf{r}^{x}$ where $\mathbf{x}_{\mathrm{sp}}$ are the fitted values (estimating the spatial part of $\mathbf{x}$) and $\mathbf{r}^{x}$ the residuals.
Estimate the covariate effect $\beta$ by fitting the model (3) but with $\mathbf{x}$ replaced by $\mathbf{r}^{x}$.
From Proposition 1, the bias in spatial+ is $\langle\mathbf{r}^{x},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{x},\mathbf{r}^{x}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$. The residuals $\mathbf{r}^{x}$ estimate the component $\mathbf{x}_{\mathrm{ns}}$ which lies in the non-spatial part of the sample space. As this subspace is orthogonal to the spatial part containing $\mathbf{z}$, this shows why we can expect the bias to be eliminated. Indeed, even though $\mathbf{r}^{x}$ may contain a small proportion of potentially confounded spatial components, its non-spatial part is likely to dominate so that bias is negligible (as seen in the Scenario 4 simulations).
@@ 5.2 Case 2: $\mathbf{x}$ is fully spatial
Sometimes $\mathbf{x}$ has no non-spatial components, i.e. $\bm{\xi}^{x}_{\mathrm{ns}}=\mathbf{0}$. For example, $\mathbf{x}$ could be entirely generated from a smooth spatial process, or the dimension $p$ of the spatial effect in (3) could be as large as the sample size $n$ (such as in a regional spatial model with one data measurement per region). The covariate vector is then contained in the space spanned by the spatial basis vectors and model (3) is unidentifiable. Here, the smoothing penalty not only improves the fit but acts as a regulariser. The model cannot be fitted without smoothing, and the apportionment of the covariate effect between the covariate and spatial terms of the model is directly related to how the smoothing is applied. Our expression for the bias still holds, however, unlike Case 1 above where the non-spatial component $\mathbf{x}_{\mathrm{ns}}$ is known to be unconfounded, we have no information of the confounding scenario and methods like spatial+ may not be effective. The behaviour of the bias will also generally be harder to predict. In particular, even if confounding is at low frequencies (making the numerator of the bias small), the bias could still be significant as the absence of non-spatial components in $\mathbf{x}$ means that the denominator could also potentially be very small, especially when the covariate is dominated by low frequency behaviours.
@@
As shown in Sections 3 and 4, spatial frequencies play a key role for the bias, and techniques such as Fourier and wavelet analysis are useful for understanding the behaviours of the observed variables $\mathbf{y}$ and $\mathbf{x}$ (as suggested, for example, in Keller and Szpiro, (2020)). However, our expressions show that it is the frequencies shared by $\mathbf{x}$ and $\mathbf{z}$ that cause bias, and the data has no information on how such frequencies are split between the two variables. Therefore, without further (untestable) assumptions for identifiability, we cannot estimate the bias.
In the case where $\mathbf{x}$ and $\mathbf{z}$ are realisations of two stationary Gaussian spatial processes, Guan et al., (2023) use a Fourier transform to describe spatial frequencies and show that an assumption of unconfoundedness at high frequencies can be used to obtain identifiability. As noted in Guan et al., (2023), this assumption seems natural as the highest frequency spatial variation in a variable is more likely to be characteristic to that particular variable and therefore not confounded with other spatial variables. Corollary 1 shows that this idea extends to our framework in which spatial frequencies are defined directly in terms of the spatial analysis model (3) as described in Section 2.2. Unconfoundedness at the highest frequencies then corresponds to the assumption that $\xi^{z}_{\mathrm{sp},i}=0$ for frequencies above a certain threshold. Thus, under this assumption, the frequency components of $\mathbf{x}$ above the threshold can be used to obtain identifiability.
We note that the frequencies used to obtain identifiability need not be the highest spatial frequencies. An example of this is the scenario of Bolin and Wallin, (2024) who consider the asymptotic behaviour of the covariate effect estimate in an analysis model of the form (3) when the true data generation process is given by
$\mathbf{y}=\beta\mathcal{S}\mathbf{x}+\bm{\epsilon},\quad\bm{\epsilon}\sim N(\bm{0},\sigma^{2}\mathbf{I})$
where $\mathcal{S}$ is a smoothing operator, for example, $\mathcal{S}\mathbf{x}$ could be the spatial frequencies of the covariate below a threshold $k$ and $\mathbf{x}-\mathcal{S}\mathbf{x}$ the remaining (higher) frequencies. Here, there is no spatial confounder as such (in the form of a missing or unmeasured spatial variable), but bias arises due to mis-specification of the covariate in the analysis model. However, since $\beta\mathcal{S}\mathbf{x}=\beta\mathbf{x}-\beta(\mathbf{x}-\mathcal{S}\mathbf{x})$, the true data generation process can be written in the form of (2) with $\mathbf{z}=-\beta(\mathbf{x}-\mathcal{S}\mathbf{x})$, and we can therefore use our framework to assess the bias in this case. From the expression for $\mathbf{z}$ we see that “confounding” is at the higher frequencies $\mathbf{x}-\mathcal{S}\mathbf{x}$ of the covariate, and the lower frequencies $\mathcal{S}\mathbf{x}$ of the covariate are unconfounded. Thus, by Corollary 1, the bias is given by $\frac{-\beta\sum_{i=k+1}^{n}(\xi^{x}_{\textrm{sp},i})^{2}w_{i}}{\sum_{i=1}^{n}(\xi^{x}_{\textrm{sp},i})^{2}w_{i}}$. As the sample size $n$ goes to infinity, we expect the higher frequencies to increasingly dominate $\mathbf{x}$ and therefore the bias to approach $-\beta$, consistent with the results of Bolin and Wallin, (2024).
@@ 5.3 Capped spatial+
Based on our analysis in Section 5.2, we present an adapted version of spatial+ which can be used to assess the bias in the more challenging Case 2 scenario. As previously noted, unconfounded high frequencies in $\mathbf{x}$ have a similar effect on the bias as non-spatial information. Therefore, assuming the $k$ highest frequencies are unconfounded, a natural way to adapt spatial+ is the following two-step procedure. Let $\mathbf{u}_{\mathrm{sp},1},\ldots,\mathbf{u}_{\mathrm{sp},n}$ denote the spatial eigenvectors in $\mathbb{R}^{n}$.
Decompose the covariate as $\mathbf{x}=\mathbf{x}_{\mathrm{sp}}^{k}+\mathbf{r}^{k}$ where $\mathbf{x}_{\mathrm{sp}}^{k}=\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ (the lowest $n-k$ frequencies) and $\mathbf{r}^{k}=\sum_{i=n-k+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ (the highest $k$ frequencies).
Estimate the covariate effect $\beta$ by fitting the model (3) but with $\mathbf{x}$ replaced by $\mathbf{r}^{k}$.
@@
Proposition 1 shows that the bias of the corresponding covariate effect estimate is $\langle\mathbf{r}^{k},\mathbf{z}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}/\langle\mathbf{r}^{k},\mathbf{r}^{k}\rangle_{\bm{\mathbf{\Sigma}}^{-1}}$, which, as $\mathbf{r}^{k}$ is assumed to be unconfounded, should therefore, in theory, be 00. However, as the spatial eigenvectors span the whole of $\mathbb{R}^{n}$, $\mathbf{r}^{k}$ and $\mathbf{z}$ no longer lie in orthogonal subspaces. As the component $\mathbf{r}^{k}$ can also be modelled as part of the highest frequencies of the spatial effect in (3), the estimate of $\beta$ obtained in this way is highly sensitive to inaccuracies at these frequencies of the estimated spatial effect. Such inaccuracies easily arise in practice as the estimation is based on overall spatial smoothing, driven by the fitted values rather than the behaviour at individual spatial frequencies. In order to avoid such sensitivity, we adapt the model matrix of the regression in (ii) by removing the columns corresponding to the frequencies above $k$. This more explicitly reflects the assumption of unconfoundedness at these frequencies but allows the spatial effect to continue to adjust for confounders at frequencies below the cap. We call the estimate of $\beta$ obtained from (i) and the adapted (ii) the capped spatial+ estimate. By construction, if the assumption of unconfoundedness above the cap is true, the bias for this estimate is 00.
In practice, the true value of the cap $k$ is unlikely to be known, and it may therefore be mis-specified at a value $k^{\prime}\neq k$. If $k^{\prime}<k$, $\mathbf{r}^{k^{\prime}}=\sum_{i=n-k^{\prime}+1}^{n}\xi^{x}_{\mathrm{sp},i}\mathbf{u}_{\mathrm{sp},i}$ is simply a smaller component of $\mathbf{r}^{k}$ which is still orthogonal to the spatial part of the adjusted spatial model. Therefore, the bias should still be eliminated. However, if $k^{\prime}>k$, then $\mathbf{r}^{k^{\prime}}$ contains a potentially confounded component of $\mathbf{x}$, namely, the $k^{\prime}-k$ highest frequencies of $\mathbf{x}_{\mathrm{sp}}^{k}$, and by Corollary 1 the bias becomes $\sum_{i=n-k^{\prime}+1}^{n-k}\xi^{x}_{\mathrm{sp},i}\xi^{z}_{\mathrm{sp},i}w_{i}/\sum_{i=n-k^{\prime}+1}^{n}(\xi^{x}_{\mathrm{sp},i})^{2}w_{i}$. For comparison, the bias in the original spatial model (3) in this scenario is $\sum_{i=1}^{n-k}\xi^{x}_{\mathrm{sp},i}\xi^{z}_{\mathrm{sp},i}w_{i}/\sum_{i=1}^{n}(\xi^{x}_{\mathrm{sp},i})^{2}w_{i}$. Thus, the numerator for the spatial model includes additional confounded frequencies contributing to the bias, but at the same time the denominator is also larger. Therefore, in this case, the capped spatial+ estimate may not reduce and could even increase the size of the bias of the spatial model.
@@
Finally, in Section 5.2 we noted that the assumption of unconfoundedness does not have to be limited to the highest frequencies of the spatial effect. If we believe the unconfounded behaviour of $\mathbf{x}$ is in a different frequency range, we can easily adapt the capped spatial+ method by replacing the component $\mathbf{r}^{k}$ by the part of $\mathbf{x}$ that is believed to be unconfounded. In the scenario of Bolin and Wallin, (2024) described at the end of Section 5.2, we could use the $k$ lowest frequencies of $\mathbf{x}$ as the component $\mathbf{r}^{k}$ under the assumption that these are contained in the unconfounded low frequency component $\mathcal{S}\mathbf{x}$ of the covariate. This is therefore also consistent with the suggestion in Bolin and Wallin, (2024) to replace $\mathbf{x}$ in the analysis model (3) by a smoothed (i.e. lower frequency) version $\hat{\mathcal{S}}\mathbf{x}$ of the covariate.